# TESSER: Transfer-Enhancing Adversarial Attacks from Vision Transformers via Spectral and Semantic Regularization

### v1

@@ 4.1 Experiment Setup
Dataset. Following prior works [38, 48, 24], we randomly selected 1,000 clean images from the ILSVRC2012 validation set [29], ensuring that all surrogate models correctly classify each image with high confidence. This selection facilitates a consistent and fair evaluation of transferability across models.
Models. We employ four representative Vision Transformer models as surrogate architectures: ViT-B/16 [6], PiT-B [17], CaiT-S24 [34], and Visformer-S [3]. To assess cross-architecture generalization, we group evaluation into two categories: ViT-to-ViT and ViT-to-CNN transfer. For ViT-to-ViT, we use four unseen target ViTs: DeiT-B [33], TNT-S [15], LeViT-256 [11], and ConViT-B [8]. For ViT-to-CNN, we evaluate against four deep CNN models: Inception-v3 (Inc-v3), Inception-v4 (Inc-v4), Inception-ResNet-v2 (IncRes-v2), and ResNet-v2-152 (Res-v2) [31, 32, 16]. Additionally, to evaluate robustness against adversarial defenses, we include three adversarially trained models: Inc-v3-ens3, Inc-v4-ens4, and IncRes-v2-adv [23, 45].
Baselines. We compare our method against a suite of strong baseline attacks. These include momentum- and variance-based methods such as MI-FGSM (MIM) [5], VMI-FGSM (VMI) [36], and Skip Gradient Method (SGM) [39]. We also include three state-of-the-art transformer-specific attacks: PNA [38], TGR [48], and ATT [24], which incorporate attention structure or token-level heuristics into their gradient manipulation strategies. Our proposed method builds on this foundation by integrating token-sensitive gradient reweighting and spectral regularization.
@@
Parameter Settings. All experiments use a maximum perturbation bound of $\epsilon=16/255$, consistent with prior work [48]. The number of PGD iterations is set to $T=10$, with a step size of $\eta=\epsilon/T=1.6/255$. Momentum is used for stabilization with decay factor $\mu=1.0$. Model- and method-specific hyperparameters follow their original settings unless otherwise stated. Input images are resized to $224\times 224$, and the patch size for transformer models is fixed at $16\times 16$. For spectral smoothness regularization, we apply Gaussian blur with fixed kernel size $(3\times 3)$ and $\sigma=0.5$. We set $\gamma_{\text{base}}=0.5$. The weakening factors $\omega$, layer truncation threshold $l_{\text{cut}}$, and the adaptive scaling factor to $\lambda$ are tuned per model to balance the influence of QKV, Attention, and MLP gradients within the backward pass. The specific values of these hyperparameters are provided in Appendix B.
@@ 4.2 Evaluating the Transferability
Ours 57.6 79.4 78.4 100.0* 75.9 83.2 85.3 69.6 78.7$\uparrow$
4.3 Ablation on Module-Wise Gradient Modulation.
@@ 4.4 Qualitative Comparison: Perturbation Semantics
To further investigate the semantic alignment of perturbations generated by our method, we visualize adversarial examples produced by ATT [24] and our proposed FSGS. Each example includes the clean image, the adversarial image, and the corresponding Grad-CAM heatmap computed from the adversarial input using the predicted (incorrect) label of the black-box model. This setup allows us to examine whether the adversarial signal still localizes on meaningful regions of the input, despite successfully inducing misclassification. As shown in Figure 2, perturbations generated by FSGS exhibit high spatial alignment with semantically salient areas (often focusing on regions corresponding to object parts or discriminative textures) even when the model’s prediction is incorrect. This behavior suggests that FSGS preserves gradients associated with tokens that encode true class-relevant features. The resulting perturbations remain semantically grounded, in contrast to ATT, which tends to spread noise across the image without semantic focus.
@@ A.2 Semantic Tokens and Gradient Alignment
This is supported by empirical findings in model interpretability [1, 20, 27] and our own Grad-CAM visualizations (see Section 4.4).
@@ A.3 Feature-Sensitive Gradient Scaling (FSGS)
[…]
@@ B.1 Algorithm
We now present the full optimization framework used to generate adversarial examples in our method. Our algorithm builds on the momentum-based PGD attack [4], and integrates three coordinated components: (1) Module and Layer-Wise Gradient Modulation to adjust the contribution of each transformer layer and suppress noisy deep-layer gradients, (2) Feature-Sensitive Gradient Scaling (FSGS) to selectively enhance semantically important token gradients, and (3) Spectral Smoothness Regularization (SSR) to constrain the perturbation’s frequency content.
@@
    Module-wise Weakening: The gradient $\mathbf{g}^{(l)}$ for each module is first scaled using a module-specific weakening factor $\omega^{(l)}\in(0,1]$ (e.g., $\omega^{(l)}_{\text{attn}},\omega^{(l)}_{\text{qkv}},\omega^{(l)}_{\text{mlp}}$). This captures prior knowledge about the sensitivity of each module.
@@
    Feature-Sensitive Gradient Scaling (FSGS): Token-level importance $\hat{\alpha}_{i}$ is computed from intermediate features, and each token’s gradient is scaled accordingly:
    $\mathbf{g}_{i}^{(l)}\leftarrow\left(\gamma_{\text{base}}+\lambda(1-\hat{\alpha}_{i})\right)\cdot\mathbf{g}_{i}^{(l)}$
@@
Input: Input $\mathbf{x}$, label $y$, model $f$, steps $T$, step size $\eta$, max perturbation $\epsilon$, Gaussian blur $\mathcal{G}_{\sigma}$, momentum factor $\mu$
@@
Initialize: $\delta^{(0)}=0$, $\mathbf{m}^{(0)}=0$ ;
@@
       $\mathbf{x}^{(t)}=\mathcal{G}_{\sigma}(\mathbf{x}+\delta^{(t-1)})$ ;
       Compute loss: $\mathcal{L}^{(t)}=\mathcal{L}(f(\mathbf{x}^{(t)}),y)$ ;
       Backpropagate with hooks at Attention, QKV, MLP ;
      foreach block $l$ do
             Extract features $\mathbf{z}^{(l)}$, gradients $\mathbf{g}^{(l)}$ ;
             Apply weakening: $\mathbf{g}^{(l)}\leftarrow\omega^{(l)}\cdot\mathbf{g}^{(l)}$ ;
             Apply layer modulation: $\mathbf{g}^{(l)}\leftarrow\tau_{l}\cdot\mathbf{g}^{(l)}$ ;
             Estimate token importance $\hat{\alpha}_{i}=\text{norm}(\mathbf{z}_{i}^{(l)})$ ;
             Apply FSGS: $\mathbf{g}_{i}^{(l)}\leftarrow(\gamma_{\text{base}}+\lambda(1-\hat{\alpha}_{i}))\cdot\mathbf{g}_{i}^{(l)}$ ;
@@
      Aggregate total gradient $\mathbf{g}^{(t)}$ and update momentum:
       $\mathbf{m}^{(t)}=\mu\cdot\mathbf{m}^{(t-1)}+\frac{\mathbf{g}^{(t)}}{\|\mathbf{g}^{(t)}\|_{1}}$ ;
      Update perturbation:
       $\delta^{(t)}=\text{Clip}_{\epsilon}(\delta^{(t-1)}+\eta\cdot\text{sign}(\mathbf{m}^{(t)}))$ ;
@@ B.2 Computational Cost
Model FSGS FSGS + SSR ATT [24]
@@ C.1 Quantitative Analysis for SSR
Table 7: The average attack success rate (%) against ViTs, CNNs, and defended CNNs by various transfer-based attacks with input diversity enhancement strategy. The best results are highlighted in bold. “PO” denotes PatchOut [38].
@@
Ours+PO 85.18$\uparrow$ 64.17$\uparrow$ 52.16$\uparrow$ Ours+PO 91.15$\uparrow$ 72.9$\uparrow$ 56.46$\uparrow$
@@
Ours+PO 94.83$\uparrow$ 87.7$\uparrow$ 61.43$\uparrow$ Ours+PO 84.42$\uparrow$ 79.4$\uparrow$ 58.06$\uparrow$
@@ C.2 Comparison of Attack Efficiency when using Input Diversity Technique
To further assess the effectiveness and generality of our proposed TESSER framework, we evaluate its performance when combined with an input diversity enhancement strategy, specifically PatchOut (PO) [38]. This technique introduces random masking during inference to improve the robustness and transferability of adversarial perturbations.
@@ Appendix D Additional Ablation Studies
To further analyze the effectiveness and interpretability of our proposed method, we present qualitative comparisons between TESSER (FSGS+SSR) and the state-of-the-art ATT [24] across a diverse set of samples from ImageNet. Figure 3 shows clean and adversarial images, Grad-CAM heatmaps, and FFT visualizations for perturbations.

### v2

@@ 4.1 Experiment Setup
Dataset. Following prior works Wei et al. (2022); Zhang et al. (2023); Ming et al. (2024), we randomly selected 1,000 clean images from the ILSVRC2012 validation set Russakovsky et al. (2015), ensuring that all surrogate models correctly classify each image with high confidence. This selection facilitates a consistent and fair evaluation of transferability across models.
Models. We employ four representative Vision Transformer models as surrogate architectures: ViT-B/16 Dosovitskiy et al. (2020), PiT-B Heo et al. (2021), CaiT-S24 Touvron et al. (2021b), and Visformer-S Chen et al. (2021). To assess cross-architecture generalization, we group evaluation into two categories: ViT-to-ViT and ViT-to-CNN transfer. For ViT-to-ViT, we use four unseen target ViTs: DeiT-B Touvron et al. (2021a), TNT-S Han et al. (2021), LeViT-256 Graham et al. (2021), and ConViT-B d’Ascoli et al. (2021). For ViT-to-CNN, we evaluate against four deep CNN models: Inception-v3 (Inc-v3), Inception-v4 (Inc-v4), Inception-ResNet-v2 (IncRes-v2), and ResNet-v2-152 (Res-v2) Szegedy et al. (2016, 2017); He et al. (2016). Additionally, to evaluate robustness against adversarial defenses, we include three adversarially trained models: Inc-v3-ens3, Inc-v4-ens4, and IncRes-v2-adv Madry et al. (2017); Xu et al. (2022).
Baselines. We compare our method against a suite of strong baseline attacks. These include momentum- and variance-based methods such as MI-FGSM (MIM) Dong et al. (2018b), VMI-FGSM (VMI) Wang and He (2021), and Skip Gradient Method (SGM) Wu et al. (2020a). We also include three state-of-the-art transformer-specific attacks: PNA Wei et al. (2022), TGR Zhang et al. (2023), and ATT Ming et al. (2024), which incorporate attention structure or token-level heuristics into their gradient manipulation strategies. Our proposed method builds on this foundation by integrating token-sensitive gradient reweighting and spectral regularization.
@@
Parameter Settings. All experiments use a maximum perturbation bound of $\epsilon=16/255$, consistent with prior work Zhang et al. (2023). The number of PGD iterations is set to $T=10$, with a step size of $\eta=\epsilon/T=1.6/255$. Momentum is used for stabilization with decay factor $\mu=1.0$. Model- and method-specific hyperparameters follow their original settings unless otherwise stated. Input images are resized to $224\times 224$, and the patch size for transformer models is fixed at $16\times 16$. For spectral smoothness regularization, we apply Gaussian blur with fixed kernel size $(3\times 3)$ and $\sigma=0.5$. We set $\gamma_{\text{base}}=0.5$. The weakening factors $\omega$, layer truncation threshold $l_{\text{cut}}$, and the adaptive scaling factor to $\lambda$ are tuned per model to balance the influence of QKV, Attention, and MLP gradients within the backward pass. The specific values of these hyperparameters are provided in Appendix B.
@@ 4.2 Evaluating the Transferability
[…]
Table 2: The attack success rate (%) of various transfer-based attacks against robust ViTs. The best results are highlighted in bold).
Model Attack Robust ViTs Normal ViTs
Swin-B Xcit-S Swin-B Xcit-S
clean 5.4 46.8 0.4 0.2
ViT-B/16 PNA+PO 8.8 51.7 47.5 45.5
TGR+PO 15.8 56.5 54.4 54.5
ATT+SPPO 16.9 56.7 70.4 68.6
TESSER 29.7 70.8 99.9 77.9
PiT-B PNA+PO 9.2 51.8 67.0 71.2
TGR+PO 17.9 58.2 77.3 80.7
ATT+SPPO 18.7 58.3 90.4 92.8
TESSER 31.9 71.6 100 95.4
We conducted additional experiments on robust ViT models trained via adversarial training with epsilon = 4, including Swin-B Mo et al. (2022) and XCiT-S Debenedetti et al. (2023). We compared TESSER against state-of-the-art attacks (PNA+PO, TGR+PO, and ATT+SPPO) using their optimal hyperparameters. As shown in Table LABEL:tab:attack_comparison_vits, TESSER consistently achieves the highest ASR on both robust and corresponding standard ViT models, confirming its strong effectiveness even under adversarial defense settings. These results demonstrate that TESSER’s transferability extends to robust ViTs, not just CNNs and hybrids.
4.3 Ablation on Module-Wise Gradient Modulation
@@ 4.4 Qualitative Comparison: Perturbation Semantics
To further investigate the semantic alignment of perturbations generated by our method, we visualize adversarial examples produced by ATT Ming et al. (2024) and our proposed FSGS. Each example includes the clean image, the adversarial image, and the corresponding Grad-CAM heatmap computed from the adversarial input using the predicted (incorrect) label of the black-box model. This setup allows us to examine whether the adversarial signal still localizes on meaningful regions of the input, despite successfully inducing misclassification. As shown in Figure 2, perturbations generated by FSGS exhibit high spatial alignment with semantically salient areas (often focusing on regions corresponding to object parts or discriminative textures) even when the model’s prediction is incorrect. This behavior suggests that FSGS preserves gradients associated with tokens that encode true class-relevant features. The resulting perturbations remain semantically grounded, in contrast to ATT, which tends to spread noise across the image without semantic focus.
@@
Evaluating TESSER Perceptual Stealthiness. While SSR reduces high-frequency components, it operates only during the perturbation generation phase. The final adversarial image is obtained by subtracting the noise and clamping the result. Thus, no direct blurring is applied to the image itself, preserving spatial clarity. To objectively assess perceptual visibility, we provide a quantitative comparison using LPIPS, SSIM, and PSNR across TESSER and transfer-based attacks such as ATT and TGR. As shown in Table 4, TESSER achieves significantly higher imperceptibility, with  50% reduction in LPIPS, 33% improvement in SSIM, and +5 dB increase in PSNR, demonstrating strong stealthiness without sacrificing effectiveness.
Table 4: Stealth Evaluation of Transfer-Based Attacks.
Metric TGR ATT TESSER Better (Stealthier)
LPIPS $\downarrow$ 0.35 0.42 0.21 TESSER (closer in perceptual similarity)
SSIM $\uparrow$ 0.66 0.57 0.77 TESSER (better structure preservation)
PSNR $\uparrow$ 22.23 dB 19.70 dB 25.04 dB TESSER (less signal distortion)
@@ A.2 Semantic Tokens and Gradient Alignment
This is supported by empirical findings in model interpretability Abnar and Zuidema (2020); Lin and Parikh (2016); Raghu et al. (2021) and our own Grad-CAM visualizations (see Section 4.4).
@@ A.3 Feature-Sensitive Gradient Scaling (FSGS)
$\cos\theta(G_{\text{FSGS}},G^{\prime})>\cos\theta(G,G^{\prime})\quad\text{(by Jensen's inequality over positively weighted aligned vectors)}$
@@ B.1 Algorithm
We now present the full optimization framework used to generate adversarial examples in our method. Our algorithm builds on the momentum-based PGD attack Dong et al. (2018a), and integrates three coordinated components: (1) Module and Layer-Wise Gradient Modulation to adjust the contribution of each transformer layer and suppress noisy deep-layer gradients, (2) Feature-Sensitive Gradient Scaling (FSGS) to selectively enhance semantically important token gradients, and (3) Spectral Smoothness Regularization (SSR) to constrain the perturbation’s frequency content.
@@
[…]
@@
    Feature-Sensitive Gradient Scaling (FSGS):
    A layer-aware gradient modulation mechanism that scales gradients based on token-wise importance. FSGS promotes perturbation alignment with semantically salient features while suppressing low-level, architecture-specific signals that degrade cross-model transferability.
    Let $\mathbf{Z}\in\mathbb{R}^{T\times D}$ be the token embedding matrix at a given transformer block. We define the raw importance score of token $i$ as: $\alpha_{i}=\|\mathbf{z}_{i}\|_{2}$. The importance scores are min-max normalized across tokens: $\hat{\alpha}_{i}=\frac{\alpha_{i}-\min_{j}\alpha_{j}}{\max_{j}\alpha_{j}-\min_{j}\alpha_{j}+\varepsilon}$, where $\varepsilon$ is a small constant to avoid division by zero.
    Let $l\in\{1,\dots,L\}$ denote the index of the current transformer block, and let $\mathcal{E}\subset\{1,\dots,L\}$ be the set of early layers (e.g., $\mathcal{E}=\{1,\dots,k\}$). Define an indicator function:
    $\beta^{(l)}=\begin{cases}1&\text{if }l\in\mathcal{E}\quad\text{(early layer)}\\ 0&\text{otherwise}\end{cases}$ (2)
    The final scaling factor for token $i$ at layer $l$ is then computed as: $s_{i}^{(l)}=\gamma_{\text{base}}+\lambda\cdot\left[(1-\beta^{(l)})\cdot\hat{\alpha}_{i}+\beta^{(l)}\cdot(1-\hat{\alpha}_{i})\right]$. And the FSGS-modulated gradient is: $\mathbf{g}_{i}^{(l),\text{FSGS}}=s_{i}^{(l)}\cdot\mathbf{g}_{i}^{(l)}$
@@
Input: Input image $\mathbf{x}$, label $y$, model $f$,
Steps $T$, step size $\eta$, perturbation bound $\epsilon$,
Gaussian blur $\mathcal{G}_{\sigma}$, momentum $\mu$,
Base scale $\gamma_{\text{base}}$,
Module-specific FSGS strengths $\lambda_{\text{qkv}},\lambda_{\text{attn}},\lambda_{\text{mlp}}$,
Early-layer set $\mathcal{E}$, attention cutoff layer $l_{\text{cut}}$,
Module weakening factors $\omega^{(l)}$,
SSR loss function $\mathcal{L}_{\text{SSR}}$
@@
Initialize: $\delta^{(0)}=0$, $\mathbf{m}^{(0)}=0$
@@
 1. Apply SSR:
 $\mathbf{x}^{(t)}=\mathcal{G}_{\sigma}(\mathbf{x}+\delta^{(t-1)})$
 2. Forward pass and compute classification loss:
 $\mathcal{L}_{\text{cls}}^{(t)}=\mathcal{L}(f(\mathbf{x}^{(t)}),y)$
 3. Backward pass with hooks at QKV, Attention, and MLP modules:
 foreach block $l\in\{1,\dots,L\}$ do
    foreach module $m\in\{\text{qkv},\text{attn},\text{mlp}\}$ do
       3.1 Extract token features and gradients:
       $\mathbf{Z}^{(l,m)}=[\mathbf{z}_{1}^{(l,m)},\dots,\mathbf{z}_{T}^{(l,m)}]$
       $\mathbf{G}^{(l,m)}=[\mathbf{g}_{1}^{(l,m)},\dots,\mathbf{g}_{T}^{(l,m)}]$
       3.2 Compute token importance:
       $\alpha_{i}=\|\mathbf{z}_{i}^{(l,m)}\|_{2}$,  $\hat{\alpha}_{i}=\frac{\alpha_{i}-\min_{j}\alpha_{j}}{\max_{j}\alpha_{j}-\min_{j}\alpha_{j}+\varepsilon}$
       3.3 Apply module-wise weakening:
       $\mathbf{G}^{(l,m)}\leftarrow\omega^{(l)}\cdot\mathbf{G}^{(l,m)}$
       3.4 Selective Attention Truncation (only if $m=\text{attn}$):
       if $l\geq l_{\text{cut}}$ then
          $\mathbf{G}^{(l,\text{attn})}\leftarrow 0$
         end if
       3.5 Compute FSGS scaling:
       foreach token $i\in\{1,\dots,T\}$ do
          if $l\in\mathcal{E}$ then
             $s_{i}=\gamma_{\text{base}}+\lambda_{m}\cdot(1-\hat{\alpha}_{i})$
            end if
          else
             $s_{i}=\gamma_{\text{base}}+\lambda_{m}\cdot\hat{\alpha}_{i}$
            end if
          $\mathbf{g}_{i}^{(l,m)}\leftarrow s_{i}\cdot\mathbf{g}_{i}^{(l,m)}$
@@ 3.5 Compute FSGS scaling:
      end foreach
   end foreach
 4. Aggregate gradients across all modules:
 $\mathbf{g}^{(t)}=\sum_{l,m}\text{Aggregate}(\mathbf{G}^{(l,m)})$
 5. Momentum update:
 6. Perturbation update with projection:
 $\delta^{(t)}=\text{Clip}_{\epsilon}(\delta^{(t-1)}+\eta\cdot\text{sign}(\mathbf{m}^{(t)}))$
@@ B.2 Computational Cost
Model FSGS FSGS + SSR ATT Ming et al. (2024)
@@ C.1 Quantitative Analysis for SSR
Table 9: The average attack success rate (%) against ViTs, CNNs, and defended CNNs by various transfer-based attacks with input diversity enhancement strategy. The best results are highlighted in bold. “PO” denotes PatchOut Wei et al. (2022).
@@
[…]
@@
[…]
@@ C.2 Comparison of Attack Efficiency when using Input Diversity Technique
To further assess the effectiveness and generality of our proposed TESSER framework, we evaluate its performance when combined with an input diversity enhancement strategy, specifically PatchOut (PO) Wei et al. (2022). This technique introduces random masking during inference to improve the robustness and transferability of adversarial perturbations.
@@ Appendix D Additional Ablation Studies
D.1 Qualitative Comparison
To further analyze the effectiveness and interpretability of our proposed method, we present qualitative comparisons between TESSER (FSGS+SSR) and the state-of-the-art ATT Ming et al. (2024) across a diverse set of samples from ImageNet. Figure 3 shows clean and adversarial images, Grad-CAM heatmaps, and FFT visualizations for perturbations.
@@ D.1 Qualitative Comparison
D.2 Impact of Module-wise Gradient Weakening
We compare ASR with and without $\omega$ (i.e., setting all $\omega$ = 1 disables gradient weakening). On ViT-B/16, using $\omega$ improves ASR from 63.0% to 83.2% (Table LABEL:tab:omega), confirming the effectiveness of selective gradient suppression.
Table 11: TESSER attack success rate (%) with and without module-wise gradient weakening $\omega$ against eight ViT models and the average attack success rate (%) of all black-box models. The best results are highlighted in bold.
ViT-B/16 w/o $\omega$ 99.6* 40.9 71 45.7 69.3 64.8 40.2 72.5 63
[…]
PiT-B w/o $\omega$ 58.1 99.9* 69.2 74.7 69 74 66.9 70.3 72.76
[…]
D.3 Impact of Selective Attention Truncation
On PiT-B, disabling attention truncation (i.e., no $l_{\text{cut}}$) leads to an average 7% drop in ASR (Table LABEL:tab:lcut), validating the importance of focusing on early-layer token gradients.
Table 12: TESSER attack success rate (%) with and without selective attention truncation $l_{\text{cut}}$ against eight ViT models and the average attack success rate (%) of all black-box models. The best results are highlighted in bold.
ViT-B/16 w/o $l_{\text{cut}}$ 100* 60.4 87.3 67.7 88.3 85.4 65.4 89.9 80.55
[…]
PiT-B w/o $l_{\text{cut}}$ 73.1 99.7* 82.1 86 84 86.4 83.2 84.1 84.82
[…]
D.4 Impact of rescaling factor
Setting $\lambda$ = 0 disables adaptive FSGS scaling (only $\gamma_{base}$ is used as a fixed multiplier). On ViT-B/16, enabling $\lambda$ improves ASR by an average of 30% (Table LABEL:tab:lambda), highlighting the value of adaptive gradient modulation in improving attack effectiveness.
In addition, we empirically validate the effectiveness of our scaling strategy by comparing it to a random scaling baseline. As shown in Table 14, our method significantly outperforms random scaling across all target models, achieving consistently higher ASR and demonstrating stronger transferability.
Table 13: TESSER attack success rate (%) with and without rescaling factor $\lambda$ against eight ViT models and the average attack success rate (%) of all black-box models. The best results are highlighted in bold.
ViT-B/16 w/o $\lambda$ 84.1* 32.1 54.5 42.3 55.1 56 43.1 56.9 53.01
[…]
PiT-B w/o $\lambda$ 51.2 92.9* 61.2 67.9 63.6 67.9 66.5 62.6 66.72
[…]
Table 14: TESSER Attack Success Rate (%) with our Scaling strategy vs. Random Scaling. Bold = better of the two scalings for the same surrogate–target pair. * denotes white-box (surrogate equals target).
Surrogate Scaling ViT-B/16 PiT-B CaiT-S/24 Visformer-S DeiT-B TNT-S LeViT-256 ConViT-B Avg
ViT-B/16 Random 86.4* 30.6 52.3 37.8 53.0 55.4 37.8 55.7 51.12
ours 100* 61.7 94.0 68.3 92.5 85.6 72.2 91.4 83.2 $\uparrow$
PiT-B Random 28.4 100* 34.6 44.8 33.9 44.1 38.3 38.2 45.28
ours 74.9 100.0* 91.6 93.2 92.1 95.0 92.4 91.7 91.4 $\uparrow$
Appendix E Evaluating the Transferability of Different Attack Methods for Targeted Attacks
While our main experiments focus on untargeted attacks, both FSGS and SSR are model-agnostic and loss-independent components applied during backpropagation. Therefore, they are fully compatible with targeted attack formulations—only the loss needs to be adapted. To validate this, we conducted targeted attack experiments using the target label set as (true label + 1). As shown in Table LABEL:tab:targeted, our method (TESSER) achieves a significantly higher targeted ASR of 43.08%, outperforming PGD (16.06%), MIM (22.01%), and ATT (33.33%), demonstrating the effectiveness and transferability of our approach in targeted settings as well. The table will be included in the revised version.
Table 15: The attack success rate (%) of various transfer-based targeted attacks against eight ViT models and the average attack success rate (%) of all black-box models. The best results are highlighted in bold.
ViT-B/16 PGD 96.1* 2.1 6.8 2.7 5.6 6 1.7 7.5 16.06
MIM 99.4* 6.6 16.1 6.3 13.9 11.8 4.4 17.6 22.01
ATT 99.5* 8.7 35.6 9.6 33.6 30.9 6.7 42.1 33.33
[…]
PiT-B PGD 0.8 96.6* 1.1 2.3 0.9 2.1 1.3 0.7 13.22
MIM 5.3 99.9* 5.1 8 4.9 6.5 4 5.5 17.4
ATT 12.1 100* 14.6 20.2 13.2 18.8 12.2 16.8 25.98
[…]
Appendix F Evaluating TESSER Performance vs. Autoattack
While AutoAttack (AA) is a strong white-box evaluation benchmark, it is not optimized for transfer-based black-box settings. To enable a fair comparison, we evaluate both TESSER and AutoAttack under the same transfer setup with a fixed perturbation budget of epsilon = 16/255. As reported in Table LABEL:tab:autoattack, TESSER achieves over 50% higher average ASR compared to AutoAttack across multiple target models. For example, when attacking PiT-B from a ViT-B/16 surrogate, TESSER achieves an ASR of 61.7%, compared to only 10.5% for AutoAttack. This gap is expected, as TESSER is explicitly designed to optimize black-box transferability, whereas AutoAttack is tailored for white-box robustness evaluation.
Table 16: The attack success rate (%) of Autoattack (AA) vs. TESSER against eight ViT models and the average attack success rate (%) of all black-box models. The best results are highlighted in bold.
ViT-B/16 AA 99.9* 10.5 42.9 13.2 33.6 38.4 17 40.4 36.98
[…]
PiT-B AA 10 98.5* 13.5 24 11.3 21.2 22.4 13.9 26.85
[…]
Appendix G Evaluation TESSER Transferability to Visual State Space Models
To further increase the architectural dissimilarity, we evaluate TESSER transferability to Vision Mamba Zhu et al. (2024)—a state-space-based architecture with bidirectional SSMs and position-aware embeddings, representing a class of models distinct from transformers. As shown in Table LABEL:tab:vim, TESSER consistently achieves the highest ASR 87.1% and 76.7% on both Vim-Tiny and Vim-Small, respectively compared to 80.9% and 69% for sota ATT attack, demonstrating robust transfer even under significant architectural and representational divergence.
Table 17: Comparative experiments of different attack methods on VIM. “clean” indicates that clean images are classified and all results indicate the percentage of classification errors (i.e., ASR).
Model Attack VIM-tiny VIM-small
ViT-B/16 clean 3.1 0.9
MIM 45.6 42
ATT 80.9 69
TESSER 87.1 76.7
PiT-B MIM 32.3 34.9
ATT 53.4 55.1
TESSER 77.4 80.7